# Strategies for implementing quantum error correction in molecular rotation

### v1

@@ Abstract
The rotation of trapped molecules offers a promising platform for quantum technologies and quantum information processing. In parallel, quantum error correction codes that can protect quantum information encoded in rotational states of a single molecule have been developed. These codes are currently an abstract concept, as no implementation strategy is yet known. Here, we present a step towards experimental implementation of these codes by introducing architecture-agnostic check and correction operators. These operators can be decomposed into elements of the quantum logic spectroscopy toolbox that is available for molecular ions. We then describe and analyze a measurement-based sequential as well as an autonomous implementation strategy in the presence of thermal background radiation, a major noise source for rotation in polar molecules. The presented strategies and methods might enable robust sensing or even fault-tolerant quantum computing using the rotation of individual molecules.
I Introduction
@@
Quantum computers promise to solve certain computational tasks more efficiently than existing classical computers, but the performance of current prototypes is limited by noise [1]. It is thus expected that quantum error correction (QEC) will be required for useful large-scale quantum information processors [2]. While the ultimate purpose of QEC protocols is to reduce errors during the execution of quantum algorithms, they can also extend the storage time of quantum memories [3]. Over the last decade, proof-of-principle QEC implementations have been demonstrated in multiple architectures [4, 5, 6, 7, 8].
@@
QEC protects quantum information by storing it redundantly, either over multiple physical information carriers, such as atomic ions or superconducting qubits [2], or in the complex Hilbert space of a single physical information carrier [9, 10, 11, 12, 13]. Here we concentrate on the latter approach, which has been used successfully in quantum harmonic oscillators. This is known as bosonic quantum error correction and has been implemented in superconducting [14] as well as trapped ion systems [15]. In the former, an extension of the storage time of a superconducting quantum memory coupled to a microwave cavity has been demonstrated [14]. Furthermore, bosonic codes can also be embedded in finite dimensional Hilbert spaces, known as qudits [9]. There, an implementation of QEC in the spin qudits of molecules has been proposed [16].
@@
Trapped molecules have recently emerged as a viable platform for quantum technologies by demonstrating entanglement [17, 18] and relatively long quantum information storage times  [19, 20]. Molecules feature multiple degrees of freedom (DOF) that can be used to store quantum information. We are interested in the rotational DOF of a single molecule which provides an infinite Hilbert space. The concept of bosonic QEC has been extended to molecular rotation in Ref. [21]. This framework has been used to develop absorption-emission codes that protect against a major source of errors in small polar molecules: spontaneous photon emission and interaction with black body radiation [22].
@@
In this work, we present and theoretically analyze implementation strategies for such codes in linear molecules using quantum logic spectroscopy (QLS) with co-trapped atomic ions [23, 24]. The Hamiltonian of the linear rotor is introduced, and absorption and emission errors in multiple regimes are discussed in section I.1. We then revisit the conditions that need to be fulfilled to correct these errors in section I.3 and summarize the construction and properties of absorption-emission codes in section I.3. Furthermore, we describe a simplified variant of the code that is able to approximately correct for the same type of errors in section I.4. We then develop architecture-agnostic check and correction operators to perform error correction that work for the original as well as the approximate codes in section II.
@@
These operators are then decomposed into elements of the quantum logic spectroscopy toolbox that is available for molecular ions. We describe the quantum logic toolbox in section III.1 and analyze a measurement-based sequential implementation strategy in section III.2. Furthermore, we develop and analyze a measurement-free autonomous implementation strategy using dissipation engineering in section III.3. We discuss the requirements that the implementation imposes on molecular species and future research directions in section IV.
I.1 The linear rotor
@@
^$\hat{\mathcal{E}}^{\textrm{BBR}}(J,m,\delta m)=\\ \sqrt{\Gamma^{\textrm{BBR}}(J,m,\delta m)}\,\hat{T}(J,m,-1,\delta m)+\textrm{h.c.}$ (7)
@@
$\Gamma^{\textrm{BBR}}(J,m,\delta m)=\\ \gamma^{\textrm{BBR}}_{J}\bigg{(}\frac{4\pi}{3}\bigg{)}|c^{J}(J-1,m+\delta m,1,\delta m)|^{2}\,,$ (8)
@@
$\mathcal{C}_{\textrm{env}}=\big{\{}\hat{E}^{\textrm{SD}}(J,\delta m),\ \hat{E}^{\textrm{BBR}}(J,\delta m)\\ :\ J\in\mathbb{N},\ \delta m\in\{-1,0,1\}\big{\}}\,.$ (11)
I.2 Error correction conditions
@@
I.3 Rotational Æ error correction codes and their properties
@@
$\mathcal{C}_{\textrm{rec}}=\big{\{}\hat{E}^{\textrm{SD}}(J_{\mathrm{C}},\delta m),\ \hat{E}^{\textrm{BBR}}(J_{\mathrm{C}},\delta m),\\ \hat{E}^{\textrm{BBR}}(J_{\mathrm{C}}+1,\delta m)\ :\ \delta m\in\{-1,0,1\}\big{\}}\,.$ (17)
@@
$\begin{split}\overline{X}&=|J_{\textrm{C}},-m_{2}\rangle\langle J_{\textrm{C}},m_{2}|+|J_{\textrm{C}},m_{1}\rangle\langle J_{\textrm{C}},-m_{1}|\\ &\quad+\textrm{h.c.}\end{split}$ (20)
@@ I.4 Approximate Æ codes
II Quantum error correction protocols
@@
In our implementation strategy, we employ spectroscopic resolvability of the $m$-sublevels for the correction of errors. In order to prevent interactions with the environment from also resolving the $m$-sublevels, we split up the correction of errors into three steps as shown in Fig. 4:
@@
Figure 4: Algorithm for sequential quantum error correction of Æ codes in a linear rotor. The $\hat{S}_{\delta J}$ check and $\delta J$-correction operations are performed repeatedly and rapidly, and the subsequent $\hat{S}_{\delta_{m}}$ check, $\delta m$-correction, and amplitude refreshment steps are performed only if a $|\delta J|=1$ error is detected. Population of the CS(7,2,5) codewords and their evolution throughout the algorithm are indicated with red and blue squares. States which are acted on by operators are highlighted.
@@
Addressing individual $m$-sublevels breaks the hierarchy of timescales for $R_{\textrm{proj}}$ in Eq. (14) and should thus only be performed if an absorption or detection event has been detected. These operations have a limited rate, but this can still be performed at $R_{\textrm{QEC}}\gg\Gamma_{\textrm{C}}$ so the rate at which a second error occurs during the correction cycle can be kept comparatively low. Additionally, these operations are required relatively seldom, namely only if an absorption or emission occurs.
@@
for $k\in\{1,2\}$. The rotation angles $\theta_{\delta J,\delta m,k}$ are determined by the $\delta J$ and $\delta m$ of the error and are derived in Appendix B.1.
@@
III Implementation strategies for trapped ions
III.1 Trapped molecular ion toolbox
@@
$\hat{H}_{\textrm{int}}({J,m,\delta J,\delta m})=\\ \hbar\Omega(J,m,\delta J,\delta m)\,\hat{T}(J,m,\delta J,\delta m)\otimes\hat{M}_{\textrm{int}}+\text{h.c.},$ (35)
@@
III.2 Sequential check operators and correction
@@
We analyze the performance of multiple correction cycles by numerical simulation using the QuTiP package [39, 40]. We choose a noise model that focuses on the inherent capabilities of the code rather than on implementation details and thus the only noise source considered is absorption and stimulated emission by BBR on the molecular rotation during the operations. While performing the $\hat{U}_{\textrm{BSB}}^{\delta J}$ check and correction operations in the unresolved regime, the interaction with BBR is modelled by solving the Lindblad master equation with collapse operators as defined in Eq. (10). This noise model preserves superpositions between the individual $m$-sublevels. On the other hand, decays which occur during $m$-resolved operations of the $m$-correction and amplitude refreshment will destroy any superposition of the Zeeman sublevels in a rotational manifold and are thus more detrimental. The decay dynamics for this resolved regime is given by the Lindblad master equation using the operators given in Eq. (7) in Appendix C. To simplify the noise model, we assume that all operations can be performed at the same timescale, which is valid if the degeneracy of the $m$-levels can be lifted with a comparatively strong AC-Stark shift as discussed in section I.1.
@@
III.3 Autonomous approximate QEC using dissipation engineering
While a sequential correction sequence which can project errors and measure their syndromes is in principle better suited to preserve coherences of the encoded logical state over long times due to the ability to refresh amplitudes, it can be difficult to implement in practice due to the large number of carefully-engineered pulses required. Here we consider a continuous approximate error correction protocol, referred to as dissipative error correction (DEC), which trades some loss of fidelity over time for less complex implementation. We first consider a scheme that extends quantum logic spectroscopy optical pumping techniques to repump molecular rotational states back to the $J_{\textrm{C}}$ manifold but without correcting errors that cause changes in $m$. Then we concatenate the implementation of a dissipative Zeeman correction scheme.
@@ III.3 Autonomous approximate QEC using dissipation engineering
where $\sqrt{4\pi/3}\,\Omega_{\downarrow,\uparrow}c^{J_{\textrm{C}}\pm 1}(J_{\textrm{C}},m,1,0)$ are the $m$-specific coupling rates and $\hat{a}^{\dagger}(\hat{a})$ are the creation(annihilation) operators for the corresponding motional mode. The $\downarrow$,$\uparrow$ , ↑ notation indicates in which direction the $\delta J$-correction is performed.
@@
$\hat{H}_{\textrm{DEC}}^{\delta\textrm{m}}=\hbar\sum_{m_{\textrm{C}}}\Big{[}\Omega_{\rightarrow}\,\hat{T}(J_{\textrm{C}},m_{\textrm{C}}-1,0,1)\otimes\hat{a}_{\rightarrow}^{\dagger}\otimes\hat{\mathbb{I}}_{\leftarrow}\\ +\Omega_{\leftarrow}\,\hat{T}(J_{\textrm{C}},m_{\textrm{C}}+1,0,-1)\otimes\hat{\mathbb{I}}_{\rightarrow}\otimes\hat{a}_{\leftarrow}^{\dagger}\Big{]}\\ +\textrm{h.c.}$ (47)
@@ IV Conclusion and outlook
Finding a molecular species that is well suited to implement the outlined strategies is a natural next step towards an experimental demonstration. Polar molecules are natural candidates as they allow $|\delta J|=1$ transitions induced by BBR and spontaneous decay, and the rotation can be controlled via electric dipole transitions. A heteronuclear diatomic molecule seems to be a good candidate as it has the fewest vibrational modes. Additionally, it has only one rotational degree of freedom. These criteria are fulfilled in a molecule with nuclear and electronic spin singlet states. Such a molecule is in the unresolved $|\delta J|=1$ regime as the transitions between different $m$-sublevels in each $J$-manifold are degenerate and thus approximates an ideal linear rotor.
@@
Our analysis shows that implementing Æ codes and error correction in the rotation of polar molecular ions seems possible. Extending these simulations to molecular systems with more complex Hamiltonians and realistic noise models for spontaneous decay and thermal radiation is required to prove their practical feasibility. Thus, future research should focus on the analysis of other sources of noise that can contribute to decoherence. Such sources of noise include magnetic and electric field fluctuations, collisions, quantum logic infidelity due to motional heating, Stark shifts, and noisy operations [44]. The full time dependence of the interaction Hamiltonians describing radiation-molecule coupling, where counter-rotating terms are included, is also relevant for some processes such as far-detuned Raman and thus is important to consider in certain systems.
@@ Acknowledgements
This research was funded by FWF 1000 Ideas project TAI-798 and ERC Horizon 2020 project ERC-2020-STG 948893. B.F. thanks Benjamin Stickler for reviewing this manuscript and discussions on system dynamics, Christopher Reilly for discussions on implementation and error models, and Christian Marciniak for discussions regarding calculations. The authors also acknowledge Victor Albert and Chin-wen Chou for discussions on Æ codes and error mitigation, Christiane Koch for discussion of coherent control, Florentin Reiter for discussions on dissipation engineering, Kenneth Brown for discussions on molecular ions, and the Quantum Optics & Spectroscopy Group and associated ion trapping groups at the Universität Innsbruck for general assistance.
@@ A.1 Rotational part of transition dipole moment
$\begin{split}c^{J_{1}}&(J_{3},m_{3},J_{2},m_{2})\\ &=\iint Y^{*}_{J_{3},m_{3}}(\theta,\phi)Y_{J_{2},m_{2}}(\theta,\phi)Y_{J_{1},m_{1}}(\theta,\phi)\\ &\quad\quad\sin{\theta}\ \mathrm{d}\theta\ \mathrm{d}\phi\\ &=\sqrt{\frac{(2J_{3}+1)(2J_{2}+1)}{4\pi(2J_{1}+1)}}C^{J_{1},0}_{J_{3},0;J_{2},0}C^{J_{1},m_{1}}_{J_{3},m_{3};J_{2},m_{2}}\\ &=\sqrt{\frac{(2J_{1}+1)(2J_{2}+1)(2J_{3}+1)}{4\pi}}(-1)^{m_{3}}\\ &\quad\quad\begin{pmatrix}J_{3}&J_{2}&J_{1}\\ -m_{3}&m_{2}&m_{1}\end{pmatrix}\begin{pmatrix}J_{3}&J_{2}&J_{1}\\ 0&0&0\end{pmatrix}\end{split}$ (48)
The Clebsch-Gordon coefficients are defined as $C^{J_{3},m_{3}}_{J_{1},m_{1};J_{2},m_{2}}=\langle J_{1},m_{1};J_{2},m_{2}|J_{3},m_{3}\rangle$.
Note that $\hat{\epsilon}\cdot\hat{r}=\sqrt{4\pi/3}\,\big{[}\epsilon_{z}Y_{1,0}(\theta,\phi)+\epsilon_{+}Y_{1,1}(\theta,\phi)+\epsilon_{-}Y_{1,-1}(\theta,\phi)\big{]}$.

### v2

@@ Abstract
The rotation of trapped molecules offers a promising platform for quantum technologies and quantum information processing. In parallel, quantum error correction codes that can protect quantum information encoded in rotational states of a single molecule have been developed. These codes are currently an abstract concept, as no implementation strategy is yet known. Here, we present a step towards experimental implementation of one family of such codes, namely absorption-emission codes. We first construct architecture-agnostic check and correction operators. These operators are then decomposed into elements of the quantum logic spectroscopy toolbox that is available for molecular ions. We then describe and analyze a measurement-based sequential as well as an autonomous implementation strategy in the presence of thermal background radiation, a major noise source for rotation in polar molecules. The presented strategies and methods might enable robust sensing or even fault-tolerant quantum computing using the rotation of individual molecules.
1 Introduction
@@
Quantum computers promise to solve certain computational tasks more efficiently than existing classical computers, but the performance of current prototypes is limited by noise [1]. It is thus expected that quantum error correction (QEC) will be required for useful large-scale quantum information processors [2]. While the ultimate purpose of QEC protocols is to reduce errors during the execution of quantum algorithms, they can also extend the storage time of quantum memories [3]. Over the last decade, proof-of-principle QEC implementations have been demonstrated in multiple architectures [4, 5, 6, 7, 8, 9], culminating most recently in several demonstrations of operation beyond break-even in different hardware architectures [10, 11].
@@
QEC protects quantum information by storing it redundantly, either over multiple physical information carriers, such as atomic ions or superconducting qubits [2], or in the complex Hilbert space of a single physical information carrier [12, 13, 14, 15, 16]. Here we concentrate on the latter approach, which has been used successfully in quantum harmonic oscillators. This is known as bosonic quantum error correction and has been implemented in superconducting [17] as well as trapped ion systems [18]. In the former, an extension of the storage time of a superconducting quantum memory coupled to a microwave cavity has been demonstrated [10, 17]. Furthermore, bosonic codes can also be embedded in finite dimensional Hilbert spaces, known as qudits [12]. There, an implementation of QEC in the spin qudits of molecules has been proposed [19, 20].
@@
Trapped molecules have recently emerged as a viable platform for quantum technologies by demonstrating entanglement [21, 22] and relatively long quantum information storage times  [23, 24]. Molecules feature multiple degrees of freedom (DOF) that can be used to store quantum information. We are interested in the rotational DOF of a single molecule which provides a large Hilbert space. The concept of bosonic QEC has been extended to molecular rotation in Ref. [25]. This framework has been used to develop absorption-emission codes that protect against a major source of errors in small polar molecules: spontaneous photon emission and interaction with black body radiation [26].
@@
In this work, we present and theoretically analyze implementation strategies for such codes in linear molecules using quantum logic spectroscopy (QLS) with co-trapped atomic ions [27, 28]. The Hamiltonian of the linear rotor is introduced, and absorption and emission errors in multiple regimes are discussed in Sec. 1.1. We then revisit the conditions that need to be fulfilled to correct these errors in Sec. 1.2 and summarize the construction and properties of absorption-emission codes in Sec. 1.3. Furthermore, we describe a simplified variant of the code that is able to approximately correct for the same type of errors in Sec. 1.4. We then develop architecture-agnostic check and correction operators to perform error correction that work for the original as well as the approximate codes in Sec. 2.
@@
These operators are then decomposed into elements of the quantum logic spectroscopy toolbox that is available for molecular ions. We describe the quantum logic toolbox in Sec. 3.1 and analyze a measurement-based sequential implementation strategy in Sec. 3.2. Furthermore, we develop and analyze a measurement-free autonomous implementation strategy using dissipation engineering in Sec. 3.3. We discuss the requirements that the implementation imposes on molecular species and future research directions in Sec. 4.
1.1 The linear rotor
@@
[…]
@@
$\Gamma^{\textrm{BBR}}(J,m,\delta m)=\gamma^{\textrm{BBR}}_{J}\bigg{(}\frac{4\pi}{3}\bigg{)}|c^{J}(J-1,m+\delta m,1,\delta m)|^{2}\,,$ (8)
@@
$\mathcal{C}_{\textrm{env}}=\big{\{}\hat{E}^{\textrm{SD}}(J,\delta m),\ \hat{E}^{\textrm{BBR}}(J,\delta m):\ J\in\mathbb{N},\ \delta m\in\{-1,0,1\}\big{\}}\,.$ (11)
1.2 Error correction conditions
@@
1.3 Rotational Æ error correction codes and their properties
@@
$\mathcal{C}_{\textrm{rec}}=\big{\{}\hat{E}^{\textrm{SD}}(J_{\mathrm{C}},\delta m),\ \hat{E}^{\textrm{BBR}}(J_{\mathrm{C}},\delta m),\hat{E}^{\textrm{BBR}}(J_{\mathrm{C}}+1,\delta m)\ :\ \delta m\in\{-1,0,1\}\big{\}}\,.$ (17)
@@
[…]
@@ 1.4 Approximate Æ codes
2 Quantum error correction protocols
@@
In our implementation strategy, we use spectroscopic resolvability of the $m$-sublevels for the correction of errors. In order to prevent interactions with the environment from also resolving the $m$-sublevels, we split up the correction of errors into three steps as shown in Fig. 4:
@@
Figure 4: Algorithm for sequential quantum error correction of Æ codes in a linear rotor. The $\hat{S}_{\delta J}$ check and $\delta J$-correction operations are performed repeatedly and rapidly, and the subsequent $\hat{S}_{\delta m}$ check, $\delta m$-correction, and amplitude refreshment steps are performed only if a ${|\delta J|=1}$ error is detected. Population of the CS(7,2,5) codewords and their evolution throughout the algorithm are indicated with red and blue squares. States which are acted on by operators are highlighted.
@@
Addressing individual $m$-sublevels breaks the hierarchy of timescales for $R_{\textrm{proj}}$ as shown in Eq. (14) and should thus only be performed if an absorption or detection event has been detected. These operations have a limited rate, but this can still be performed at ${R_{\textrm{QEC}}\gg\Gamma_{\textrm{C}}}$ so the rate at which a second error occurs during the correction cycle can be kept comparatively low. Additionally, these operations are required relatively seldom, namely only if an absorption or emission occurs.
@@
for ${k\in\{1,2\}}$. The rotation angles $\theta_{\delta J,\delta m,k}$ are determined by the $\delta J$ and $\delta m$ of the error and are derived in Appendix B.
@@
3 Implementation strategies for trapped ions
3.1 Trapped molecular ion toolbox
@@
[…]
@@
3.2 Sequential check operators and correction
@@
We analyze the performance of multiple correction cycles by numerical simulation using the QuTiP package [43, 44]. We choose a noise model that focuses on the inherent capabilities of the code rather than on implementation details and thus the only noise source considered is absorption and stimulated emission by BBR on the molecular rotation during the operations. While performing the $\hat{U}_{\textrm{BSB}}^{\delta J}$ check and correction operations in the unresolved regime, the interaction with BBR is modelled by solving the Lindblad master equation with collapse operators as defined in Eq. (10). This noise model preserves superpositions between the individual $m$-sublevels. On the other hand, decays which occur during $m$-resolved operations of the $m$-correction and amplitude refreshment will destroy any superposition of the Zeeman sublevels in a rotational manifold and are thus more detrimental. The decay dynamics for this resolved regime is given by the Lindblad master equation using the operators given in Eq. (7) in Appendix C. To simplify the noise model, we assume that all operations can be performed at the same timescale, which is valid if the degeneracy of the $m$-levels can be lifted with a comparatively strong AC-Stark shift as discussed in Sec. 1.1.
@@
3.3 Autonomous approximate QEC using dissipation engineering
While a sequential correction sequence which can project errors and measure their syndromes is in principle better suited to preserve coherences of the encoded logical state over long times due to the ability to refresh amplitudes, it can be difficult to implement in practice due to the large number of carefully-engineered pulses required. Here we consider a continuous approximate error correction protocol, referred to as dissipative error correction (DEC), which trades some loss of fidelity over time for less complex implementation. Conditions for autonomous error correction similar to the Knill-Laflamme conditions have been found in Ref. [45]. Here we focus on approximate schemes, considering first a scheme that repumps molecular rotational states back to the $J_{\textrm{C}}$ manifold but without correcting errors that cause changes in $m$. Then we concatenate the implementation of a dissipative Zeeman correction scheme.
@@ 3.3 Autonomous approximate QEC using dissipation engineering
[…]
@@
[…]
@@ 4 Conclusion and outlook
Finding a molecular species that is well suited to implement the outlined strategies is a natural next step towards an experimental demonstration. Polar molecules are natural candidates as they allow ${|\delta J|=1}$ transitions induced by BBR and spontaneous decay, and the rotation can be controlled via electric dipole transitions. A heteronuclear diatomic molecule seems to be a good candidate as it has the fewest vibrational modes. Additionally, it has only one rotational degree of freedom. The simplest system which can host Æ codes in rotational DOF without any other contributions to angular momentum is a molecule with nuclear and electronic spin singlet states. Such a molecule is in the unresolved ${|\delta J|=1}$ regime as the transitions between different $m$-sublevels in each $J$-manifold are degenerate and thus approximates an ideal linear rotor.
@@
Our analysis shows that implementing Æ codes and error correction in the rotation of polar molecular ions seems possible. Extending these simulations to molecular systems with more complex Hamiltonians and realistic noise models for spontaneous decay and thermal radiation is required to prove their practical feasibility. Thus, future research should focus on the analysis of other sources of noise that can contribute to decoherence. Such sources of noise include magnetic and electric field fluctuations, collisions, quantum logic infidelity due to motional heating and logic ion spontaneous decay, Stark shifts, and noisy operations [49]. The full time dependence of the interaction Hamiltonians describing radiation-molecule coupling, where counter-rotating terms are included, is also relevant for some processes such as far-detuned Raman and thus is important to consider in certain systems. Experimental milestones in the degree of quantum control needed to realize such QEC codes have recently been achieved in molecules and atoms. State preparation, state-resolved readout, and coherent control of CaH+ in ${J\leq 3}$, utilizing QLS via a co-trapped Ca+ logic ion with microwave carrier and CW Raman sideband transitions has been demonstrated [50].
@@
The Æ codes considered here can be cast in terms of dephasing codes, discussed in Jain, et al. [26] Recent theoretical work has extended these code families to more efficient encodings and has uncovered links to spin codes and binomial codes [53]. Investigation of the feasibility and performance of these dephasing codes against realistic dephasing noise, as well as hypothetical collision codes which protect against angular momentum transfers due to collisions with background gas, will be explored in future work.
@@ Acknowledgements
This research was funded by FWF 1000 Ideas project TAI-798 and ERC Horizon 2020 project ERC-2020-STG 948893. The authors thank Benjamin Stickler for reviewing the manuscript and discussions on system dynamics, Christopher Reilly for discussions on implementation and error models, and Christian Marciniak for discussions regarding calculations. The authors also acknowledge Victor Albert and Chin-wen Chou for discussions on Æ codes and error mitigation, Christiane Koch for discussion of coherent control, Florentin Reiter for discussions on dissipation engineering, Kenneth Brown for discussions on molecular ions, and the Quantum Optics & Spectroscopy Group and associated ion trapping groups at the Universität Innsbruck for general assistance.
@@
Data Availability
The source code for the simulations that support the findings of this study are openly available at https://doi.org/10.5281/zenodo.14536088.
@@ A.1 Rotational part of transition dipole moment
$\begin{split}c^{J_{1}}(J_{3},m_{3},J_{2},m_{2})&=\iint Y^{*}_{J_{3},m_{3}}(\theta,\phi)Y_{J_{2},m_{2}}(\theta,\phi)Y_{J_{1},m_{1}}(\theta,\phi)\ \sin{\theta}\ \mathrm{d}\theta\ \mathrm{d}\phi\\ &=\sqrt{\frac{(2J_{3}+1)(2J_{2}+1)}{4\pi(2J_{1}+1)}}C^{J_{1},0}_{J_{3},0;J_{2},0}C^{J_{1},m_{1}}_{J_{3},m_{3};J_{2},m_{2}}\\ &=\sqrt{\frac{(2J_{1}+1)(2J_{2}+1)(2J_{3}+1)}{4\pi}}(-1)^{m_{3}}\\ &\quad\quad\begin{pmatrix}J_{3}&J_{2}&J_{1}\\ -m_{3}&m_{2}&m_{1}\end{pmatrix}\begin{pmatrix}J_{3}&J_{2}&J_{1}\\ 0&0&0\end{pmatrix}\end{split}$ (48)
The Clebsch-Gordon coefficients are defined as ${C^{J_{3},m_{3}}_{J_{1},m_{1};J_{2},m_{2}}=\langle J_{1},m_{1};J_{2},m_{2}|J_{3},m_{3}\rangle}$ and it should be noted that $\hat{\epsilon}\cdot\hat{r}=\sqrt{4\pi/3}\,\big{[}\epsilon_{z}Y_{1,0}(\theta,\phi)+\epsilon_{+}Y_{1,1}(\theta,\phi)+\epsilon_{-}Y_{1,-1}(\theta,\phi)\big{]}$.